Solve for x the following quadratic equations by factoring
X squared - 5x +6=0

Answer by TimothyLamb(4379)

xx - 5x + 6 = 0
(x - 2)(x - 3) = 0
---
two real roots at:
x = 2
x = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php